OpenStudy (anonymous):

Okay 4 sqrt 10 times 2 sqrt 90 is??

OpenStudy (anonymous):

4sqrt(10)*2sqrt(90)=4sqrt(10)*2sqrt(9*10)=4sqrt(10)*6sqrt(10)=sqrt(10)[4+6]=10sqrt(10)

OpenStudy (radar):

\[4\sqrt{10}\times2\sqrt{90}=4\sqrt{10}\times2\sqrt{9\times10}=4\sqrt{10}\times6\sqrt{10}\]

OpenStudy (radar):

Now you can multiply getting \[24\sqrt{10}\]

OpenStudy (anonymous):

Thanks!